• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 08:40:54 on April19,2001

    Entry number 60787

    keyword=compton dipole, bogies left

    Accelerator reported at the 8:00 am meeting that the compton dipole still had 1A of current in it when it was supposed to be at 0 current. [That has the potential to cause steering problems for beam into the hall.] Accelerator has asked AES to fix it so that when 0A is requested from the box supply, 0A is delivered.

    The left bogies may have a new "feature". Last night and again this morning, after clicking on "PSM enable", "DM enable", "brake release" (and getting yes back from each), when clicking on "clamp release" and getting yes, it immediately turns to no when the mouse is moved away. Shift crews should keep an eye on this and report if it continues to happen.